OpenStudy (anonymous):

Carbon-14 has a half-life of 5,730 years. In how many years will 120 grams of Carbon-14 decay to 15 grams? 11,460 years 17,190 years 22, 920 years 28,650 years

OpenStudy (anonymous):

@doafan could you help :)

OpenStudy (anonymous):

B. You need to divide 120 by 2 three times to get to 15 grams. 120 ÷ 2 = 60 (1 time) 60 ÷ 2= 30 (2 times) 30 ÷ 2= 15 (3 times) So the C14 (Carbon-14) decayed until only 15 grams are left. Now, using the amount of years for the half life (5,730) multiply it by the number of times it has decayed, which is 3 times. 5730 × 3 = 17,190

OpenStudy (anonymous):

thank you soooo much!!!

OpenStudy (anonymous):

@livenlearn no problem!! :D